OpenStudy (anonymous):

you have $50,000 and wish to invest part at 10% and the rest at 6%. How much should be invested at each rate to produce the same return as if it all had been invested at 9%?

OpenStudy (amistre64):

a + b = 50 ; a = 50 - b ................................... a(.1)+ b(.06) = 50(.09) (50-b)(.1) +b(.06) = 4.5 5 - b(.1) + b(.06) = 4.5 b(.06 - .1) = 4.5 - 5 b = -.5/-.04 = 50/4 = 12.5

OpenStudy (amistre64):

Does: 37,500 (.1) + 12,500(.06) = 50,000(.09) ?